Estimate the value of the equilibrium constant at 635 K for each of the following reactions. ?G?f for BrCl(g) is ?1.0 kJ/mol. The standard molar entropy, S?, for BrCl(g) is 240.0 J/mol?K.

2NO2(g)?N2O4(g). ?H?f for N2O4(g) is 9.16 kJ/mol.

Br2(g)+Cl2(g)?2BrCl(g) .?H?f for BrCl(g) is 14.6 kJ/mol.
Answer #1

From text book

compound AGp(kJ/mol)
NO2 51.3
N2O4 99.8

2NO2(g)<===> N2O4(g).

\Delta G_r^{\circ} =\sum n\times \Delta G^{\circ}(products)-\sum n\times \Delta G^{\circ}(reactants)

▲Go-99.8-2 × 51.3

\Delta G_r^{\circ} =-2.8 \ kJ/mol

Formula to find k

\Delta G_r^{\circ} =-RTlnK

K =e^{\frac{-\Delta G^{\circ}}{RT}}

K =e^{\frac{-(-2.8\times 10^3 J/mol)}{635 K\times 8.314 J/K/mol}}

K = 1.70

----------------------------------------

Br2(g)+Cl2(g)<===>2BrCl(g)

BrCl(g) is 14.6 kJ/mol.

compound \Delta H_f^{\circ}(kJ/mol) AGp(kJ/mol) S_f^{\circ}(J/mol/K)
BrCl(g) 14.6 -1.0 240.0
Br2(g) 30.9 3.1 245.5
Cl2(g) 0 0 223.1

\Delta G^{\circ}= 2(-1.0)-3.1 =-5.1\ kJ

At 635 K

\Delta G^{\circ}= -RTlnK

K=e^{\frac{-\Delta G}{RT}}

K=e^{\frac{-(-5.1\times 10^{-3})}{8.314\times 635}}

K = 2.63